OpenStudy (anonymous):

The width of a rectangle is 7 less than its length. If the area of the rectangle is 198, find the perimeter of the rectangle.

OpenStudy (anonymous):

P = 2L + 2W L = x W = x - 7...the width is 7 less than the length

OpenStudy (anonymous):

Now, you put the variables in for L and W P = 2(x) + 2(x - 7)

OpenStudy (anonymous):

P = 2x+ 2x - 14 P = 4x - 14

OpenStudy (anonymous):

l = 7+w A = l x w 198 = w(7 + w) w^2 + 7w - 198 = 0 (w + 18)(w - 11) = 0 w = 11 because it cant be negative l = 18 P = 2(w + l) = 2(11 + 18) = 58 so perimeter = 58m